MONT 106N – Identifying Patterns Discussion Day 2 October 2, 2009

Josh Hauser and Patrick Jarvis

In class on October 2, 2009, you will be responsible for about 20 minutes of the class period, during which we will discuss the article “Two Mammals’ Longevity Boosted” from the Wall Street Journal, published on July 9, 2009. You can assume everyone will read the article before class – the information how to get it is posted on the course homepage. So you don’t need to spend a lot of time summarizing what it says. Instead, come up with several good questions that you can use to start the discussion, and then be prepared to ”moderate” (maybe answer some questions if people have not understood some point, call on people who have something to say, and so forth).

Some suggestions: Try to focus about half of the time on discussing this technical point: The article reports some numbers about average life spans and says a 14% increase “is really extreme.” How would a scientist come to that conclusion? To really understand whether that is a big increase, what additional information would we also need to know? (Idea is: If we did the experiment multiple times, the averages of the lifespans of the treatment and control groups would also have some histograms describing their distributions, so averages and SD’s. The comparison is being made based on that idea. The implicit claim is that an increase as big as 14% is a big number of SD’s over the average lifespans of control groups – something we will discuss in detail later in the course! Be sure to talk to me before the discussion if you want to go over this.) Then, open things up and ask: What practical implications would there be if human life span could be increased by these methods? Should we be trying to extend human life spans this way? What should a non-expert know about stuff like this to be well-informed?

I found this by looking at a listing in Chance News 52 – see link to this on the course homepage. You can also get to the original version of the article from there. You will see discussion of the issues in Chance News that may give you a few ideas of how to get a discussion started. And of course, you can talk things over with me before class too.